OpenStudy (anonymous):

Fun Land theme park has a total of 80 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? 80 60 30 20

OpenStudy (anonymous):

i need help i think its 60

terenzreignz (terenzreignz):

lol Let's pretend there aren't any choices for the moment :) First step to such problems is to represent the values: If we let x = number of males employed How would you represent the number of females?

OpenStudy (anonymous):

y?

hero (hero):

Hint: m + f = 80 m = 2f + 20

terenzreignz (terenzreignz):

That approach would make sense, in this setup, as you know that there are ONLY 80 employees in all, you know that if there are x males, the rest must be female, how many are the rest? :)

OpenStudy (anonymous):

m+f=80 males= 20 more than twice the number of females substitute m=2f+20 into m+f=80 2f+20+f=80 Find the number for females employed and you can find the number of males employed using the m=2f+20 equation.

OpenStudy (anonymous):

i have another question....What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14

OpenStudy (anonymous):

Hint: 2y=4 y=2

terenzreignz (terenzreignz):

It seems easier to get the y value first. Subtract the second equation from the first, you'll get the equation @monokerous has provided.

OpenStudy (anonymous):

i got that one can u help with the last one i put

OpenStudy (anonymous):

you need to substitute y=2 into either one of your equations to find x.

OpenStudy (anonymous):

ok but im still confused how do u solve it

terenzreignz (terenzreignz):

Well, as mentioned, you substitute y=2 to any of the two equations like so: 4x + 3(2) = 10 Now solve for x :)

OpenStudy (anonymous):

o yay i get it because 4+6=10 and 4+10=14 duh

OpenStudy (anonymous):

o forgot :(

OpenStudy (anonymous):

i think its one right

terenzreignz (terenzreignz):

Isn't it strange that simply removing the x made the equations true? then x can only be... :D

terenzreignz (terenzreignz):

lol, yes, you got it :D

OpenStudy (anonymous):

can u help me wif another one

terenzreignz (terenzreignz):

Sure, I'll do what I can ^.^

OpenStudy (anonymous):

thanks What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18 2 1 -2 -1

terenzreignz (terenzreignz):

Can't guarantee that the answers will look as nice though :) Are you willing to work with a bit of decimals?

OpenStudy (anonymous):

would it be 2 because 13*2=26 - 6=20 and 7*2=14+4=18 so y would equal 1 again

terenzreignz (terenzreignz):

Hang on, let's do this step by step, because I'm a bit confused (sorry :( ) First, multiply the first equation by 2, on both sides And multiply the second equation by 3, on both sides.

terenzreignz (terenzreignz):

You were right about y=1 though. If you're confident about your solution, then you should be fine :)

OpenStudy (anonymous):

ok thanks ill try its just a pretest but im actually trying to get a good grade lol